# Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model

### v1

@@ Abstract
Introduction.– In cold atom systems with strong interactions [RevModPhys.80.885, Kinoshita2006, Lewenstein2007, Gross2017, Zhang2018] and disorder [Garreau2017], dissipation is an ever-present adversary, since environmental coupling tends to destroy intriguing, but delicate, quantum phenomena. However, it is being increasingly appreciated that dissipation itself may in fact be harnessed [diehl2008quantum, Bardyn_2013, goldman2016topological, Müller2021, Helmrich2020, Sponselee_2019, Gou_2020, Liang_2022, Bouchoule2021, Yamamoto2023, mengdissipative, fandissipative, Gong2018, ashida2020non, bergholtzreview] and induce unique intriguing effects [Gong2018, ashida2020non, bergholtzreview], thus providing a new frontier in the study of non-equilibrium quantum many-body systems [eisert2015quantum]. The vast majority of recent studies have focused on non-interacting dissipative systems [Gong2018, ashida2020non, bergholtzreview]. In the quantum realm these systems may realize novel topological phenomena such as the Liouvillian version [Song2019, PhysRevLett.127.070402, PhysRevResearch.4.023160] of the non-Hermitian skin effect [lee2016, yao2018, kunst2018, martinez2018, Okuma2020, Okuma2023, lin2023] in which a macroscopic number of eigenstates are exponentially localized at the boundaries of the system. Several recent studies investigate the interplay between interactions and the skin effect [Yoshida2023, Kim2023, Yang2021, Mao_2023, Wang_2023, zheng2023exact, fate22, Shen_2022, hamanaka2024multifractality, zhong2024density], yet this topic remains far less explored than the non-interacting case.
@@
Figure 1: The models under consideration in this letter. (a) A triangular ladder as implemented in Refs. Gou_2020 and Liang_2022, with an additional on-site potential and interactions of strength $U$. The upper sites have dissipation with rate $\Gamma$, while the lower ones do not. A fictitious magnetic flux $\phi$ leads to a relative phase in the hopping terms between the different sites Upon integrating out the green sites we obtain (b) with dissipation acting on neighbouring sites, where the magnetic flux has been chosen to be $\phi=-\frac{\pi}{2}$. (c) Shows the system described by the integrable unidirectional effective non-Hermitian (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model), where the dissipation has been absorbed into the hopping parameter and the on-site potential.
@@
$H=-$ $t\sum_{j=1}^{N}a_{j}^{\dagger}a_{j+1}+a_{j+1}^{\dagger}a_{j}+\frac{U}{2}\sum_{j=1}^{N}n_{j}(n_{j}-1)-\sum_{j=1}^{N}\mu_{j}n_{j},$
@@
where $a_{j}$ are bosonic annihilation operators, $n_{j}=a^{\dagger}_{j}a_{j}$ are the number operators, and $\mu_{j}$ are random disorder variables.
One way to incorporate realistic dissipation so that the system becomes exactly solvable is to consider a triangular ladder [Liang_2022, Gou_2020], as in Fig. 1(a). Particles dissipate from the sites on the top and these sites can be integrated out [supplemental_material] (see also Ref. Kamenevbook, leaving a chain with jump operators
$\ell_{j}=a_{j}+ia_{j+1}$ (3)
@@
$H_{eff}$ $=-2t\sum_{j=1}^{N}a_{j+1}^{\dagger}a_{j}+\frac{U}{2}\sum_{j=1}^{N}n_{j}(n_{j}-1)-\sum_{j=1}^{N}\mu_{j}^{\prime}n_{j}$
$-\epsilon a^{\dagger}_{1}a_{N}.$ (4)
where we defined $\mu^{\prime}_{j}=\mu_{j}-2it$ for $j\neq 1,N$ and $\mu^{\prime}_{1}=\mu_{1}-i(t+\frac{\epsilon}{2}),\mu^{\prime}_{N}=\mu_{N}-i(t+\frac{\epsilon}{2})$. Ref. Liang_2022 realized and analyzed the non-interacting and non-disordered limit of Eq. (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model) in a cold atom system. Here we include both interactions and disorder, and in addition to Eq. (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model), also consider the full Liovillian quantum dynamics as described by Eqs. (1), (2) and (3).
@@
Correlation functions and phase diagram for open boundary conditions.– We now examine the properties of the model defined by the effective Hamiltonian (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model) with random disorder and open boundary conditions ($\epsilon=0$), which exhibits a remarkable degree of solvability. As we detail below this leads to the phase diagram in Fig. 2. For arbitrary disorder the Bethe equations now take the factorized form
@@
Figure 2: The ground state phase diagram of the non-Hermitian Hamiltonian with open boundary conditions at integer filling. For $U>2\Delta$ the ground state is in a Mott phase, which can either be localized to one of the boundaries or to the right depending on $t$, whereas for $U<2\Delta$ the system is in Bose-Glass-like phase with several different condensates, which may also be localized to either the left or the right. At $U=0$ there is a transition between a Bose-Einstein condensate at the boundary and a Bose-Einstein condensate at a single site in the bulk.
@@
which are eigenstates if $\lambda=\mu_{n}^{\prime}$ for $n=1,2,\ldots N$. If there is no disorder, meaning $\mu_{n}^{\prime}=0,$ all the eigenstates coalesce to $\psi(x)=\delta_{x,1}$. One simple observation is in that if $\lambda=\mu_{n}^{\prime}$ then the particle density is 00 for $x>m$ where $m$ is the smallest integer such that $\mu_{m}^{\prime}=\mu_{n}^{\prime}$. The particle is therefore sharply localized to a finite region. In addition, if $t>>|\mu_{m^{\prime}}^{\prime}-\mu_{n^{\prime}}^{\prime}|$ for all $m^{\prime},n^{\prime}$ we see that the particle is exponentially localized in the left side of that region for all eigenstates. On the other hand, if $t<<|\mu_{m^{\prime}}^{\prime}-\mu_{n^{\prime}}^{\prime}|$ on average each eigenstate is exponentially localized to its corresponding site. Extending to the many-body ground state, we see that it is a gapless condensate of particles in the state with $\lambda$ equal to the $\mu_{n}^{\prime}$ with smallest real part. Due to the simplicity of this state, it is possible to obtain the asymptotics of the two-point correlation function. Let $x_{2}>x_{1}$ both be in the region where the wave function is non-zero. Then, using the notation $\mu_{k}=\mu_{k}^{\prime}+2it$
@@
$\bra{GS}a^{\dagger}_{x_{1}}a_{x_{2}}\ket{GS}\sim\prod_{n=1}^{x_{1}}\left(\frac{\mu_{n}-\mu_{x}}{-2t}\right)^{2}\prod_{n=x_{1}+1}^{x_{2}}\left(\frac{\mu_{n}-\mu_{x}}{-2t}\right).$ (13)
In order to proceed we specify that the disorder variables $\mu_{k}$ are uniformly distributed in the interval $[-\Delta,\Delta]$. Since this is the ground state $\mu_{n}$ must be the smallest disorder variable, which is $-\Delta$ in the thermodynamic limit. Including the wave function normalization and averaging over the disorder we find that the asymptotics of the two point correlator is
$G(x_{1},x_{2})$ $\propto\left(\frac{\Delta}{-t}\right)^{2x_{1}}\left(\frac{\Delta}{-2t}\right)^{x_{2}-x_{1}}.$ (14)
@@
As expected this is not translationally invariant, and we see that the correlation length $\xi=1/\ln\frac{\Delta}{2t}$ which diverges at $\Delta=2t,$ signifying a phase transition. One of the phases is a skin-effect-phase, where all the particles are localized at one of the edges (corresponding to the red line at the left edge of Fig. 2), and the other phase is a disorder-induced localized phase, where all the particles are localized at some site (corresponding to the green line at the left edge of Fig. 2). Other states show a similar behaviour.
@@
Now we let $U>0$. This has dramatic effects for arbitrarily weak interactions in the thermodynamic limit, $N\rightarrow\infty$. In the ground state the particles are no longer condensed into the same mode, due to the repulsive interactions. The particles spread out over the different modes, so that there are competing condensates. This is reminiscent of a Bose-Glass [Fisher1989, Scalettar1991], where the competition between disorder and interactions causes disjoint localized pockets of superfluid condensate to form. Here, in contrast, the condensates are not spatially disjoint unless the disorder is much larger than the hopping parameter, as in the non-interacting case.
@@
If $U$ is sufficiently small one of the modes will still have a most of the particles, say $M_{0}$, so that there are rapidities $\lambda_{n}=-\Delta+(n-1)U$ for $n=1,2,\ldots M_{0}$. The magnitude of the disorder part in the position dependent factor in the wave function is always smaller than $2\Delta,$ so a factor $\left(\frac{(M_{0}-1)U+2\Delta}{-2t}\right)^{x_{1}+x_{2}}$ may be extracted from the sum, so that for $x_{2}>>x_{1}$ the leading behaviour of the two-point correlator is approximately
$G(x_{1},x_{2})\sim\left(C\frac{(M_{0}-1)U+2\Delta}{-2t}\right)^{x_{2}-x_{1}}$ (15)
@@
for some real number $C\sim 1$ which arises from the disorder average. We see that the correlation length is $\xi=1/\ln\left(\frac{(M_{0}-1)U+2\Delta}{2t}\right).$ Its divergence at $2t=2\Delta+(M_{0}-1)U$ marks a transition between a boundary-localized (the purple part of Fig. 2) and a site-localized (the green part of Fig. 2) phase.
@@
If $U>2\Delta$ the energy is minimized by equally distributing the integers in Eq. (11). Suppose for simplicity that $M/N=n$ is an integer. In that case the gap is $U-2\Delta$, so this ground state is a Mott insulator at integer filling. The leading order behaviour of the two-point correlator for $x_{2}>>x_{1}$ is
$G(x_{1},x_{2})\sim\left(C\frac{(n-1)U+2\Delta}{-2t}\right)^{x_{2}-x_{1}}$ (16)
again for some real number $C\sim 1$. Note that at $n=1$ the leading order behaviour is independent of $U$ (up to the constant $C,$ which may have some $U$ dependence). There is a phase transition at $(n-1)U+2\Delta=\frac{2t}{C}$ between a site-localized and a boundary-localized Mott insulator, as shown in the yellow and red part of Fig. 2 respectively.
@@
Liouvillian dynamics.– Now let us turn to the dynamics of the full Lindblad master equation. For a generic initial state we expect that the time evolution includes every eigenoperator (9), hence the two-point correlator consists of some time-dependent linear combination of all two-point correlators of the model defined by the Hamiltonian (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model). By the same reasoning as for the ground state correlators the leading order time dependent asymptotics are
$G(x_{1},x_{2};\tau)=\sum_{n=1}^{M}D_{n}(\tau)\left(C_{n}\frac{(n-1)U+2\Delta}{-2t}\right)^{x_{2}-x_{1}},$ (17)
@@
where $\tau$ is the time, $C_{n}\sim 1$ are constants, and $D_{n}(\tau)$ are obtained from the coefficients in Eq. (9). Since there is only dissipation it follows that the functions $D_{n}(\tau)\rightarrow 0$ as $\tau\rightarrow\infty$. We see that if $U$ is non-zero there will only be a phase transition between boundary localization and site localization once the mean particle number has decreased to a finite number. If $U=0,$ on the other hand, there is always such a dynamic phase transition for $t=\Delta$.
@@
As mentioned above, the realization of a momentum-space lattice version of this model in the non-interacting limit in an cold atom setup was reported in Ref. Gou_2020. There the parameters were chosen so that interactions were unimportant and attractive. We expect that similar setups can be tuned to the regime of repulsive interactions described in this paper, although the real-space interactions must be chosen to be attractive in order the be repulsive in the momentum-space lattice [An_2018]. While previous realizations of the non-interacting limit have used ${}^{87}\mathrm{Rb}$ atoms, ${}^{39}\mathrm{K}$ may be more suitable for the interacting system, as its relevant Feschbach resonances are wider, allowing for more precise control of the interactions [Meierthesis].
@@
Finally, we also note that an alternative interpretation of the non-Hermitian Hamiltonian Eq.  (Liouvillian skin effects and fragmented condensates in an integrable dissipative Bose-Hubbard model) can be gleaned from [Hatano1996, Hatano1997, Lehrer1998]. There the difference in the hopping to the left and to the right, respectively, are obtained from the strength of a magnetic field applied to interacting flux lines in an array of defects in a superconductor. The unidirectional Hamiltonian is obtained when the transverse field component is large.

### v3

@@ Abstract
Introduction.– In cold atom systems with strong interactions [RevModPhys.80.885, Kinoshita2006, Lewenstein2007, Gross2017, Zhang2018] and disorder [Garreau2017], dissipation is an ever-present adversary, since environmental coupling tends to destroy intriguing, but delicate, quantum phenomena. However, it is being increasingly appreciated that dissipation itself may in fact be harnessed [diehl2008quantum, Bardyn_2013, goldman2016topological, Müller2021, Helmrich2020, Sponselee_2019, Gou_2020, Liang_2022, Bouchoule2021, Yamamoto2023, mengdissipative, fandissipative, Gong2018, ashida2020non, bergholtzreview] and induce unique intriguing effects [Gong2018, ashida2020non, bergholtzreview, grun2022protocol], thus providing a new frontier in the study of non-equilibrium quantum many-body systems [eisert2015quantum]. The vast majority of recent studies have focused on non-interacting dissipative systems [Gong2018, ashida2020non, bergholtzreview]. In the quantum realm these systems may realize novel topological phenomena such as the Liouvillian version [Song2019, PhysRevLett.127.070402, PhysRevResearch.4.023160, Mao_2024] of the non-Hermitian skin effect [lee2016, yao2018, kunst2018, martinez2018, Okuma2020, Okuma2023, lin2023] in which a macroscopic number of eigenstates are exponentially localized at the boundaries of the system. Several recent studies investigate the interplay between interactions and the skin effect [Yoshida2023, Kim2023, Yang2021, Mao_2023, Wang_2023, zheng2023exact, fate22, Shen_2022, hamanaka2024multifractality, zhong2024density], yet this topic remains far less explored than the non-interacting case.
@@
Figure 1: The models under consideration in this letter. (a) A triangular ladder as implemented in Refs. Gou_2020 and Liang_2022, with an additional on-site potential and interactions of strength $U$. The upper sites have dissipation with rate $\Gamma$, while the lower ones do not. A fictitious magnetic flux $\phi$ leads to a relative phase in the hopping terms between the different sites. Upon integrating out the green sites we obtain (b) with dissipation acting on neighbouring sites, where the magnetic flux has been chosen to be $\phi=\frac{\pi}{2}$. (c) Shows the system described by the integrable unidirectional effective non-Hermitian (4), where the dissipation has been absorbed into the hopping parameter and the on-site potential.
@@
$H=-$ $t\sum_{j=1}^{N}a_{j}^{\dagger}a_{j+1}+a_{j+1}^{\dagger}a_{j}+\frac{U}{2}\sum_{j=1}^{N}n_{j}(n_{j}-1)+\sum_{j=1}^{N}\mu_{j}n_{j},$
@@
where $a_{j}$ are bosonic annihilation operators, $n_{j}=a^{\dagger}_{j}a_{j}$ are the number operators, and $\mu_{j}$ are disorder variables.
One way to incorporate realistic dissipation so that the system becomes exactly solvable is to consider a triangular ladder [Liang_2022, Gou_2020], as in Fig. 1(a). Particles dissipate from the sites on the top and these sites can be integrated out [supplemental_material] (see also Ref. Kamenevbook), leaving a chain with jump operators
$\ell_{j}=a_{j}-ia_{j+1}$ (3)
@@
$H_{eff}\!$ $=\!-2t\sum_{j=1}^{N}a_{j}^{\dagger}a_{j+1}\!+\!\frac{U}{2}\sum_{j=1}^{N}n_{j}(n_{j}-1)\!+\!\sum_{j=1}^{N}\mu_{j}^{\prime}n_{j}\!-\!\epsilon a^{\dagger}_{1}a_{N}.$ (4)
where we defined $\mu^{\prime}_{j}=\mu_{j}-2it$ for $j\neq 1,N$ and $\mu^{\prime}_{1}=\mu_{1}-i(t+\frac{\epsilon}{2}),\mu^{\prime}_{N}=\mu_{N}-i(t+\frac{\epsilon}{2})$. Ref. Liang_2022 realized and analyzed the non-interacting and non-disordered limit of Eq. (4) in a cold atom system. Here we include both interactions and disorder, and in addition to Eq. (4), also consider the full Liovillian quantum dynamics as described by Eqs. (1), (2) and (3). An immediate conclusion that can be drawn is that in the absence of interactions a state with $N$ particles decays at a rate of $-\mathrm{Im}\left(\sum_{j}\mu^{\prime}_{j}n_{j}\right)\approx 2tN$.
@@
Correlation functions and phase diagram for open boundary conditions.– We now examine the properties of the model defined by the effective Hamiltonian (4) with random disorder and open boundary conditions ($\epsilon=0$), which exhibits a remarkable degree of solvability. As we detail below this leads to the phase diagram in Fig. 2. For arbitrary disorder the Bethe equations now take the factorized form
@@
Figure 2: The ground state phase diagram of the non-Hermitian Hamiltonian with open boundary conditions at integer filling as obtained analytically from the two-point correlation functions in Eqs. (15) and (16). For $U>2\Delta$ the ground state is in a Mott phase, which can either be localized at the boundary or be evenly distributed depending on $t$, whereas for $U<2\Delta$ the system is in Bose-Glass-like phase of fragmented condensates, which may also be localized either at the boundary or at some site. At $U=0$ there is a transition between a Bose-Einstein condensate at the boundary and a Bose-Einstein condensate at a single site in the bulk.
@@
which are eigenstates if $\lambda=\mu_{n}^{\prime}$ for $n=1,2,\ldots N$. If there is no disorder, meaning $\mu_{n}^{\prime}=0$ for all $n,$ all the eigenstates coalesce to $\psi(x)=\delta_{x,1}$. One simple observation is in that if $\lambda=\mu_{n}^{\prime}$ then the particle density is 00 for $x>m$ where $m$ is the smallest integer such that $\mu_{m}^{\prime}=\mu_{n}^{\prime}$. The particle is therefore sharply localized to a finite region. In addition, if $t>>|\mu_{m^{\prime}}^{\prime}-\mu_{n^{\prime}}^{\prime}|$ for all $m^{\prime},n^{\prime}$ we see that the particle is exponentially localized in the left side of that region for all eigenstates. On the other hand, if $t<<|\mu_{m^{\prime}}^{\prime}-\mu_{n^{\prime}}^{\prime}|$ on average each eigenstate is exponentially localized to its corresponding site. Extending to the many-body ground state, we see that it is a gapless condensate of particles in the state with $\lambda$ equal to the $\mu_{n}^{\prime}$ with smallest real part. Due to the simplicity of this state, it is possible to place strong bounds on the two-point correlation function. Let $x_{2}>x_{1}$ both be in the region where the wave function is non-zero. Note that the imaginary parts of the $\mu_{n}$:s cancel in the eigenstates, except for the factors that include $\mu_{1}^{\prime}$ and $\mu_{N}^{\prime}$. These do not have a qualitative effect on the physics, so we neglect them. Using the notation $\mu_{k}=\mu_{k}^{\prime}+2it$
@@
$G(x_{1},x_{2})$ $=\frac{\bra{GS}a^{\dagger}_{x_{1}}a_{x_{2}}\ket{GS}}{\langle GS|GS\rangle}\!\propto\prod_{j=1}^{x_{1}}\!\left(\frac{\mu_{n}\!-\!\mu_{j}}{-2t}\right)\!\prod_{j=1}^{x_{2}}\left(\frac{\mu_{n}\!-\!\mu_{j}}{-2t}\right).$ (13)
In order to proceed we must place some constraints on the disorder variables. Suppose that they are bounded and not all the same. Then we may choose the smallest interval $\mathcal{I}$ such that $\mu_{j}\in\mathcal{I}$ for all $j$ to be $\mathcal{I}=[-\Delta,\Delta]$ with $\Delta>0$. Clearly $\mu_{n}=-\Delta$. Suppose further that the least $\mu_{m}>-\Delta$ is given by $\mu_{m}=-\Delta+\delta$. Then the two-point correlation function satisfies the bounds
$\frac{1}{\langle GS|GS\rangle}\left(\frac{\delta}{2t}\right)^{x_{1}+x_{2}}\leq|G(x_{1},x_{2})|\leq\frac{1}{\langle GS|GS\rangle}\left(\frac{\Delta}{t}\right)^{x_{1}+x_{2}}$ (14)
@@
for $x_{1},x_{2}<n$. If either $x_{1}$ or $x_{2}$ is greater than or equal to $n$ then $G(x_{1},x_{2})=0$. We see that the correlation length satisfies $1/\xi\in\!\left[\ln\frac{\delta}{2t},\ln\frac{\Delta}{t}\right]$ and that $\xi$ diverges for some $t\leq\Delta$ signifying a phase transition. One of the phases is a skin-effect-phase, where all the particles are localized at one of the edges (corresponding to the red line at the left edge of Fig. 2), and the other phase is a disorder-induced localized phase, where all the particles are localized at some site (corresponding to the green line at the left edge of Fig. 2). Other states show a similar behaviour.
@@
Now we let $U>0$. This has dramatic effects for arbitrarily weak interactions in the thermodynamic limit, $N,M\rightarrow\infty$. In the ground state the particles no longer condense into the same mode, due to the repulsive interactions. Instead, the particles spread out over the different modes, so that there are competing condensates. The Bose-Einstein condensate has fragmented due to the interactions. This is reminiscent of a Bose-Glass [Fisher1989, Scalettar1991], where the competition between disorder and interactions causes disjoint localized pockets of superfluid condensate to form. Here, in contrast, the condensates are not spatially disjoint unless the disorder is much larger than the hopping parameter.
@@
If $U$ is sufficiently small one of the modes will still have the largest fraction of particles, say $M_{0}>M/N$ of them, so that there are rapidities $\lambda_{n}=-\Delta+(n-1)U$ for $n=1,2,\ldots M_{0}$. The leading part of the two-point correlator can then be shown to satisfy
$D\!\left(\frac{\!(M_{0}\!-\!1)U\!+\!\delta}{2t}\right)^{X}\!\leq$ $|G(X)|\leq\!D\!\left(\frac{\!(M_{0}\!-\!1)U\!+\!2\Delta}{2t}\right)^{X}$ (15)
@@
where we defined $X=x_{1}+x_{2}$ and $D>0$. We see that the correlation length satisfies $1/\xi\in\!\left[\ln\frac{(M_{0}-1)U+\delta}{2t},\ln\frac{(M_{0}-1)U+2\Delta}{2t}\right]$. Clearly $\xi$ diverges at some point, marking a transition between a boundary-localized skin-effect phase (the purple part of Fig. 2) and a site-localized phase (the green part of Fig. 2) where $M_{0}$ particles are localized at the site where the disorder potential is minimized, while the rest are distributed in condensates elsewhere in the lattice.
@@
If $U>2\Delta$ the energy is minimized by equally distributing the integers in Eq. (11). Suppose for simplicity that $M/N=n$ is an integer. In that case the gap is $U-2\Delta$, so this ground state is a Mott insulator at integer filling. The leading order behaviour of the two-point correlator satisfies
$D\left(\frac{(n-1)U+\delta}{2t}\right)^{X}\!\leq$ $|G(X)|\!\leq D\left(\frac{(n-1)U+2\Delta}{2t}\right)^{X}$ (16)
again for $D>0$. The correlation length satisfies $1/\xi\in\!\left[\ln\frac{(n-1)U+\delta}{2t},\ln\frac{(n-1)U+2\Delta}{2t}\right]$ and diverges at some point, signaling a phase transition between a boundary-localized Mott-skin effect phase and a Mott phase where the particles are evenly distributed across the lattice, as shown in the yellow and red parts of Fig. 2, respectively.
Numerical calculations corroborating these results are presented in the supplemental material [supplemental_material].
@@
Liouvillian dynamics.– Now let us turn to the dynamics of the full Lindblad master equation. For a generic initial state we expect that the time evolution includes every eigenoperator (9), hence the two-point correlator consists of some time-dependent linear combination of all two-point correlators of the model defined by the Hamiltonian (4). By the same reasoning as for the ground state correlators the leading order time dependent asymptotics at time $\tau$ have the form $G(X;\tau)=\sum_{n=1}^{M}D_{n}(\tau)g_{n}(X)$ where
$\left(\frac{(n-1)U+\delta}{2t}\right)^{X}\!\leq|g_{n}(X)|\!\leq\left(\frac{(n-1)U+2\Delta}{2t}\right)^{X}$ (17)
@@
and the $D_{n}(\tau)$ are obtained from the coefficients in Eq. (9). Since there is only dissipation it follows that the functions $D_{n}(\tau)\rightarrow 0$ as $\tau\rightarrow\infty$. We see that if $U$ is non-zero there will only be a phase transition between boundary localization and site localization once the mean particle number has decreased to a finite number. If $U=0,$ on the other hand, there is always such a dynamic phase transition.
@@
As mentioned above, the realization of a momentum-space lattice version of this model in the non-interacting limit in a cold atom setup was reported in Ref. Gou_2020. There the parameters were chosen so that interactions were unimportant and attractive. We expect that similar setups can be tuned to the regime of repulsive interactions described in this paper, although the real-space interactions must be chosen to be attractive in order the be repulsive in the momentum-space lattice [An_2018]. While previous realizations of the non-interacting limit have used ${}^{87}\mathrm{Rb}$ atoms, ${}^{39}\mathrm{K}$ may be more suitable for the interacting system, as its relevant Feschbach resonances are wider, allowing for more precise control of the interactions [Meierthesis]. The existing experimental implementations of the non-interacting limit can therefore be modified to include interactions, which we expect would exhibit the phase transitions described above.
@@
Finally, we also note that an alternative interpretation of the non-Hermitian Hamiltonian Eq.  (4) can be gleaned from [Hatano1996, Hatano1997, Lehrer1998]. There the difference in the hopping to the left and to the right, respectively, are obtained from the strength of a magnetic field applied to interacting flux lines in an array of defects in a superconductor. The unidirectional Hamiltonian is obtained when the transverse field component is large.